OpenStudy (anonymous):

angela's average for 6 math tests is 87. on her first four test she had scores of 90 , 87, 82 , and 89. on her last test , she scored 14 points lower than she did on her fifth test. what scores did angela receive on her 5th and 6th test?

OpenStudy (agreene):

(90+87+82+89+x+[x-14])/6 = 87 you get this because the arithmetic mean is the sum of the parts divided by how many parts, and it tells you it = 87. Then we just solve... start by combining like pairs and get: (334+2x)/6 = 87 multiply both sides by 6 334+2x = 522 subtract your 334 2x = 188 divide by 2 x = 94 now we know on the 5th test, she got a 94. and since the 6th test is 14 less... 94-14 = 80 on her 6th test.

OpenStudy (anonymous):

the sum of her 4 math test scores is: 87*6=522 so 522=90+87+82+89+x+x-14 solve the equation